OpenStudy (anonymous):

Gain a medal! Find the horizontal or oblique asymptote of f(x) = -2x^2+3x+6/x+1 y=-1 y=-2 y=2x+3 y=-2+5

OpenStudy (anonymous):

from the graph there are no horizontal asymptotes and there are no oblique asymptotes. there is a horizontal and parabolic asymptotes but no vertical or oblique asymptotes

OpenStudy (anonymous):

which means what? Algebra really isn't my strong suit :p @habibmatatta

OpenStudy (anonymous):

it means that the answer is that "there are no vertical or oblique asymptotes". thats the answer

OpenStudy (anonymous):

That's not an answer choice.

OpenStudy (anonymous):

plus, it was asking for horizontal; not vertical.

OpenStudy (anonymous):

sorry there isn't a horizontal there is a vertical asymptote

OpenStudy (ybarrap):

-2 x^2+3 x+6 = (5-2 x) × (x+1)+1 So, f(x) = (-2 x^2+3 x+6) / (x+1) = (5-2 x) + 1/(x+1) This equation has an oblique asymptote: 5 - 2x, because as x goes to infiity, the last term goes to zero. As x goes to -1, f(x) goes to infinity, so there is a vertical asymptote at x=-1.

OpenStudy (anonymous):

But @ybarrap she is looking for horizontal or oblique asymptotes

OpenStudy (ybarrap):

There is no horizontal asymptote, just vertical and oblique.

OpenStudy (anonymous):

whats the oblique one. is it -2x^2+3x+1? thats what i got

OpenStudy (ybarrap):

5 - 2x is the oblique one